Evaluate each of the following limits. Simplify all answers completely. Show all work toward your answer or provide an explanation for how you arrived at your answer.


1) lim x--> (x^2 -1)/(x^2 -8x+7)


2) lim x-->0^negative(-) (2)/(tan(x))


3) lim x--> infinity (cos(t))/(e^3t)


4) lim x--> (-)infinity (-3x^3 +5x^2 -4)




1
Expert's answer
2021-09-27T15:08:57-0400

1)


"\lim\limits_{x\to1}\dfrac{x^2-1}{x^2-8x+7}=\lim\limits_{x\to1}\dfrac{(x-1)(x+1)}{(x-1)(x-7)}"

"=\lim\limits_{x\to1}\dfrac{x+1}{x-7}=\dfrac{1+1}{1-7}=-\dfrac{1}{3}"

2)


"\lim\limits_{x\to0^-}\dfrac{2}{\tan x}=-\infin"

"=\lim\limits_{x\to1}\dfrac{x+1}{x-7}=\dfrac{1+1}{1-7}=-\dfrac{1}{3}"



3)


"\lim\limits_{t\to\infin}\dfrac{\cos t}{e^t}=0"

4)


"\lim\limits_{x\to-\infin}(-3x^3+5x^2-4)=\infin"